# CUSIDE-T: Chunking, Simulating Future and Decoding for Transducer based Streaming ASR

### v1

@@ 1 Introduction
Streaming ASR (also known as online ASR) is very important for many real-world applications, whose goal is to emit recognition results as quickly and accurately as possible on the fly when the user is speaking. Recently, streaming methods have been mainly studied for end-to-end ASR. There exist three end-to-end (E2E) approaches: connectionist temporal classification (CTC) [1], recurrent neural network transducer (RNN-T) [2], and attention-based encoder-decoder (AED) [3, 4, 5]. While the common for all these approaches is to define a differentiable sequence-level loss of mapping the acoustic sequence to the label sequence, the three approaches do have different properties, especially for streaming ASR.
@@ 2.2 Chunk-based training
With the steaming-friendly nature of RNN-T, streaming and non-streaming modes only differ in the acoustic encoder networks. Conventional uni-directional recurrent neural networks like LSTM [14] are able to perform forward pass in a strictly streaming mode. However, they are not widely used in recent end-to-end ASR systems due to their higher error rates [15, 16], compared to attention-based encoders like Transformer [17] and Conformer [16]. It is challenging to support streaming in attention-based models, which benefit from the full-context modeling. To make a trade-off between recognition accuracy and streaming latency, many previous works use chunk-based training instead [8, 11, 10, 18, 19], i.e., restricting the attention layers to capture a limited range of context (the chunk). In implementation, chunking-based training can be simply done via masking the attention matrix or splicing input utterances to chunking clips.
@@ 3 CUSIDE-T
In this section, we give details of our design of CUSIDE-T, which incorporats CUSIDE within RNN-T for streaming ASR and uses external language model fusion. Overview of CUSIDE-T is shown in Figure 1.
@@ 4.1 Datasets
In our experiments, the CUSIDE-T models are trained on two Chinese Mandarin labeled speech datasets respectively: the AISHELL-1 dataset [12] and the WenetSpeech dataset [13]. The AISHELL-1 contains a 150-hour training set, a 20 hour development set, and a 10-hour test set. The test set contains 7176 utterances in total. WentSpeech is a large dataset of industrial scale in fields such as entertainment, journalism, literature, technology and free conversation and so on. It is collected from online websites YouTube and Podcast and transcribed through optical character recognition (OCR) and ASR and contains a 10,000-hour training set, and we use $dev$ as development set and $test\_net$ , $test\_meeting$ that contains about 33,100 utterances in total as the test set for evaluation. Further, a benchmarking dataset SpeechIO [^1] https://github.com/SpeechColab/Leaderboard is used for additional evaluation. The SpeechIO testsets consist of 18 datasets SpeechIO-{01-18} collected from various scenarios.
@@ 4.3 AISHELL-1 Task
To compare CUSIDE-T and U2++, we first evaluate on the AISHELL-1 dataset. The released code from U2++ is used, where dynamic chunk-based attention strategy is used in training. For decoding with U2++, we test with two settings for the chunk size, 400ms and 640ms, respectively. For CUSIDE-T, left context size is 800ms for history information, and right context size is 400ms for future information. During training, the chunk size is randomly sampled from $\text{Unif}(C-A,C+A)$. We experiment with with two settings for $C$, 400ms and 640ms, respectively, both with $A$=200ms. Speed perturbation with 0.9, 1.0, 1.1 is applied on the whole data.
@@ 5 Conclusion
In this paper, we propose CUSIDE-T, which integrates the recent CUSIDE methodology into the RNN-T model for streaming ASR. We present the use of language model rescoring to further improve the performance. Experiments on both AISHELL-1 and the larger WenetSpeech dataset clearly show that the streaming ASR performance of CUSIDE-T is better than U2++. This confirms the advantage of future context simulation in the CUSIDE methodology. This work also reveals the superiority of LM rescoring over attention rescoring. The code, scripts and checkpoints used in this paper will be open source for reproduction upon the acceptance of this work.

### v2

@@ 1 Introduction
Streaming ASR (also known as online ASR) is very important for many real-world applications, whose goal is to emit recognition results as quickly and accurately as possible on the fly when the user is speaking. Recently, streaming methods have been mainly studied for end-to-end ASR. There exist three end-to-end (E2E) approaches: connectionist temporal classification (CTC) [1], recurrent neural network transducer (RNN-T) [2], and attention-based encoder-decoder (AED) [3, 4, 5]. While it is common for all these approaches to define a differentiable sequence-level loss of mapping the acoustic sequence to the label sequence, the three approaches do have different properties, especially for streaming ASR.
@@ 2.2 Chunk-based training
With the streaming-friendly nature of RNN-T, streaming and non-streaming modes only differ in the acoustic encoder networks. Conventional uni-directional recurrent neural networks like LSTM [14] are able to perform forward pass in a strictly streaming mode. However, they are not widely used in recent end-to-end ASR systems due to their higher error rates [15, 16], compared to attention-based encoders like Transformer [17] and Conformer [16]. It is challenging to support streaming in attention-based models, which benefit from the full-context modeling. To make a trade-off between recognition accuracy and streaming latency, many previous works use chunk-based training instead [8, 11, 10, 18, 19], i.e., restricting the attention layers to capture a limited range of context (the chunk). In implementation, chunking-based training can be simply done via masking the attention matrix or splicing input utterances to chunking clips.
@@ 3 CUSIDE-T
In this section, we give details of our design of CUSIDE-T, which incorporates CUSIDE within RNN-T for streaming ASR and uses external language model fusion. Overview of CUSIDE-T is shown in Figure 1.
@@ 4.1 Datasets
In our experiments, the CUSIDE-T models are trained on two Chinese Mandarin labeled speech datasets respectively: the AISHELL-1 dataset [12] and the WenetSpeech dataset [13]. The AISHELL-1 contains a 150-hour training set, a 20 hour development set, and a 10-hour test set. The test set contains 7176 utterances in total. WenetSpeech is a large dataset of industrial scale in fields such as entertainment, journalism, literature, technology and free conversation and so on. It is collected from online websites YouTube and Podcast and transcribed through optical character recognition (OCR) and ASR and contains a 10,000-hour training set, and we use $dev$ as development set and $test\_net$ , $test\_meeting$ that contains about 33,100 utterances in total as the test set for evaluation. Further, a benchmarking dataset SpeechIO [^1] https://github.com/SpeechColab/Leaderboard is used for additional evaluation. The SpeechIO testsets consist of 18 datasets SpeechIO-{01-18} collected from various scenarios.
@@ 4.3 AISHELL-1 Task
To compare CUSIDE-T and U2++, we first evaluate on the AISHELL-1 dataset. The released code from U2++ is used, where dynamic chunk-based attention strategy is used in training. For decoding with U2++, we test with two settings for the chunk size, 400ms and 640ms, respectively. For CUSIDE-T, left context size is 800ms for history information, and right context size is 400ms for future information. During training, the chunk size is randomly sampled from $\text{Unif}(C-A,C+A)$. We experiment with two settings for $C$, 400ms and 640ms, respectively, both with $A$=200ms. Speed perturbation with 0.9, 1.0, 1.1 is applied on the whole data.
@@ 5 Conclusion
In this paper, we propose CUSIDE-T, which integrates the recent CUSIDE methodology into the RNN-T model for streaming ASR. We present the use of language model rescoring to further improve the performance. Experiments on both AISHELL-1 and the larger WenetSpeech dataset clearly show that the streaming ASR performance of CUSIDE-T is better than U2++. This confirms the advantage of future context simulation in the CUSIDE methodology. This work also reveals the superiority of LM rescoring over attention rescoring. The code, scripts and checkpoints used in this paper are available at https://github.com/thu-spmi/CAT.